# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v2

@@ D.2. Upper Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
Proposition D.7 follows from the definition of a segment tree $\mathcal{T}_{N}$ and canonical partition $\text{\sf CP}_{N}$. For example, consider the segment tree $\mathcal{T}_{8}$ depicted in Figure 2. Let’s take the segment $[2,5]$. Its canonical partition consists of $\{[2,2],[3,4],[5,5]\}$. Note that only $[2,2]$ and $[5,5]$ are singleton intervals. But $2$ and $5$ are the endpoints of the interval $[2,5]$. The above proposition says that we cannot find in the canonical partition of any segment $[\alpha,\beta]$ any singleton intervals that are not end points of $[\alpha,\beta]$.
@@
Theorem D.8 ().
[…]
@@
The proof is very similar to the proof of Theorem 3.4 from Section D.1 for $\text{IVM}^{\pm}[Q]$. For every insert or delete $\delta_{\tau}\mathcal{D}=\{sR_{j}(\bm{t})\}$ where $s\in\{+,-\}$, we now have to insert an extra tuple $+\overline{R}_{j}([\tau,\tau],\bm{t})$ into $\overline{\mathcal{D}}$. This only increases the size of $\overline{\mathcal{D}}$ by a constant factor. Hence, it doesn’t affect the overall runtime.
@@
The containment $\subseteq$ in (74) holds thanks to the extra intervals$+\overline{R}_{j}([\tau,\tau],\bm{t})$ that we are inserting to $\overline{\mathcal{D}}$ in the delta version of Algorithm 1. The other containment $\supseteq$ in (74) follows from Proposition D.7. In particular, (73) only selects the leaf corresponding to $[\tau,\tau]$ in the segment tree, but Proposition D.7 tells us that this leaf could not belong to the canonical partition of any interval $[\alpha,\beta]$ unless $\tau\in\{\alpha,\beta\}$. Such intervals could only have been added to $\overline{\mathcal{D}}$ during the update $\delta_{\tau}\mathcal{D}$ because the timestamp $\tau$ is unique to this update.
@@
D.3. Upper bound in terms of $|\mathcal{D}|$ instead of $N$
@@
[…]
@@
(79) $\sum_{i\in[k-1]}|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})}$
(80) $=$ $\sum_{i\in[k-1]}\sum_{\tau\in[\tau_{i},\tau_{i+1})}\frac{1}{\tau_{i+1}-\tau_{i}}\cdot|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})}$
@@
D.4. Lower Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
@@
For the rest of this section, we fix a query $Q$, its time extension $\overline{Q}$ (Definition 6.2), and one component $\widehat{Q}_{\bm{\sigma}}$ in its multivariate extension $\widehat{Q}$ (Definition 2.3). Moreover, we fix an arbitrary database instance $\widehat{\mathcal{D}}_{\bm{\sigma}}$ for $\widehat{Q}_{\bm{\sigma}}$. Algorithm 2 describes our reduction from $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Theorem C.9 already gives a reduction from $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ to the (static) evaluation of $\overline{Q}$ over some database instance $\overline{\mathcal{D}}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$ that is given by Definition C.7 and satisfies the following properties:
@@
(Recall the definition of $H_{k}$ from Theorem C.9.) The above properties follow from Proposition C.8 and Theorem C.9, which also imply that $\overline{\mathcal{D}}$ can be constructed in linear time in $|\widehat{\mathcal{D}}_{\bm{\sigma}}|$, and $H_{k}(\overline{Q}(\overline{\mathcal{D}}))$ can be computed in linear time in $|\overline{Q}(\overline{\mathcal{D}})|=|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|$. It remains to show how to reduce the (static) evaluation of $\overline{Q}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q]$. This is shown by the following Lemma.
@@
  • $\overline{\mathcal{D}}\leftarrow\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$triangleright Definition C.7
@@
  • return $\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})=H_{k}(\overline{Q}(\overline{\mathcal{D}}))$triangleright Theorem C.9
Lemma D.9 ().
[…]
Assuming Lemma D.9 holds, Theorem 3.6 follows by taking $f(N)=N^{\omega(\widehat{Q}_{\bm{\sigma}})-1-\gamma}$ for some constant $\gamma>0$. (Recall Definition 2.2.) By Lemma D.9, this would lead into an algorithm for $\text{Eval}[\overline{Q}]$ with runtime $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\omega(\widehat{Q}_{\bm{\sigma}})-\gamma}+|\overline{Q}(\overline{\mathcal{D}})|)$, and by extension, an algorithm for $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ with runtime $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{\bm{\sigma}}|^{\omega(\widehat{Q}_{\bm{\sigma}})-\gamma}+|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|)$. This would contradict the definition of $\omega(\widehat{Q}_{\bm{\sigma}})$, thus proving that our choice of $f(N)$ is not possible. Next, we prove Lemma D.9.
Proof of Lemma D.9.
@@
Appendix E Missing Details in Section 3
@@
Proposition E.5 ().
[…]
E.1. Proof of Proposition E.1
Proposition E.1.
@@
It follows from Proposition E.2 that the multivariate extension $\widehat{Q}$ of any non-hierarchical query $Q$ is not acyclic. This is because acyclic queries have a fractional hypertree width of $1$. It remains to prove:
Lemma E.6 ().
@@
We are ready to prove Lemma E.6. Consider a hierarchical query $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$ and its multivariate extension $\widehat{Q}=\bigvee_{\bm{\sigma}\in\Sigma_{k}}\widehat{Q}_{\bm{\sigma}}$ (Eq (3)). We show that each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ is acyclic. WLOG we only consider the component for $\bm{\sigma}=(1,\ldots,k)$:
@@
This query has $k-1$ atoms and contains the variables $Z_{2},\ldots,Z_{k}$ but not $Z_{1}$. Observe that $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ is the multivariate extension of a query $Q^{\prime}$ that results from $Q$ by skipping the atom $R_{1}(\bm{X}_{1})$. Hierarchical queries remain hierarchical under atom elimination. Hence, $Q^{\prime}$ is hierarchical. By induction hypothesis, $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ is acyclic. Since we obtained $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ from $\widehat{Q}_{\bm{\sigma}}$ by eliminating root variables, leaf variables, or subsumed atoms, it follows from Claim E.1 that $\widehat{Q}_{\bm{\sigma}}$ must be acyclic. This completes the induction step.
E.2. Proof of Proposition E.2
Proposition E.2.
@@
Therefore, $\rho^{*}(\widehat{Q}_{\bm{\sigma}}^{\prime})=3/2$. We use this fact to to show that $\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$ must be at least $3/2$. To that end, we use variable elimination orders (see Appendix A). In particular, we use the definition of fractional hypertree width in terms of variable elimination orders, as given by Eq. (16). Let $\bm{\mu}$ be an optimal variable elimination order for $\widehat{Q}_{\bm{\sigma}}$, i.e., $\mathsf{w}(\bm{\mu})=\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$. We make the following claim:
Claim E.2 ().
@@
The first equality above follows from our assumption that $\bm{\mu}$ is an optimal variable elimination order for $\widehat{Q}_{\bm{\sigma}}$. The second inequality above follows from the monotonicity of the fractional edge cover number $\rho^{*}$ (see Appendix A). ∎
E.3. Proof of Proposition E.3
Proposition E.3.
@@
Consider an arbitrary Loomis-Whitney query $Q$ of some degree $k\geq 3$ and its multivariate extension $\widehat{Q}$ with fresh variables $Z_{1},\ldots,Z_{k}$. Proposition E.2 implies that the fractional hypertree width of $\widehat{Q}$ is at least $\frac{3}{2}$. Let $\widehat{Q}_{\bm{\sigma}}$ be an arbitrary component of $\widehat{Q}$. Next, we show that the fractional hypertree width of $\widehat{Q}_{\bm{\sigma}}$ is at most $\frac{3}{2}$, which concludes the proof.
Let $\widehat{Q}_{\bm{\sigma}}^{\prime}$ be the query that we obtain from $\widehat{Q}_{\bm{\sigma}}$ by eliminating the variable $Z_{k}$. Since this variable occurs in only a single atom in $\widehat{Q}_{\bm{\sigma}}$, the fractional hypertree width of $\widehat{Q}_{\bm{\sigma}}$ and $\widehat{Q}_{\bm{\sigma}}^{\prime}$ must be the same. This follows from the equivalence between variable elimination orders and tree decompositions, as explained in Section A. It remains to show that $\widehat{Q}_{\bm{\sigma}}^{\prime}$ has a fractional hypertree width of $\frac{3}{2}$. By definition of a multivariate extension, $\widehat{Q}_{\bm{\sigma}}^{\prime}$ must contain three atoms of the form:
@@
E.4. Proof of Proposition E.4
Proposition E.4.
@@
E.5. Proof of Proposition E.5
Proposition E.5.
[…]
The proof is based on the Online Vector-Matrix-Vector Multiplication (OuMv) Conjecture, which is implied by the Online Matrix-Vector Multiplication (OMv) Conjecture (Henzinger et al., 2015). See Appendix A.
Consider the simplest non-hierarchical query $Q_{nh}=R(A)\wedge S(A,B)\wedge T(B)$. The proof of Proposition E.5 consists of three parts. First, we show hardness for the problem $\text{IVM}^{\pm}[Q_{nh}]$ (Section E.5.1). Afterwards, we show how the hardness proof for $\text{IVM}^{\pm}[Q_{nh}]$ extends to a hardness proof for $\text{IVM}^{\pm}_{\delta}[Q_{nh}]$ (Section E.5.2). Finally, we explain how these two results generalize to arbitrary non-hierarchical queries without self-joins (Section E.5.3). The technique used in the latter part is standard in the literature (see, e.g. (Berkholz et al., 2017)), so we only give the high-level idea.
E.5.1. Hardness of $\text{IVM}^{\pm}[Q_{\text{nh}}]$

### v3

@@ E.2. Upper Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
Proposition E.7 follows from the definition of a segment tree $\mathcal{T}_{N}$ and canonical partition $\text{\sf CP}_{N}$. For example, consider the segment tree $\mathcal{T}_{8}$ depicted in Figure 10. Let’s take the segment $[2,5]$. Its canonical partition consists of $\{[2,2],[3,4],[5,5]\}$. Note that only $[2,2]$ and $[5,5]$ are singleton intervals. But $2$ and $5$ are the endpoints of the interval $[2,5]$. The above proposition says that we cannot find in the canonical partition of any segment $[\alpha,\beta]$ any singleton intervals that are not end points of $[\alpha,\beta]$.
@@
Theorem E.8.
The delta version of Algorithm 2 solves the $\text{IVM}^{\pm}_{\delta}[Q]$ with amortized update time $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}-1})$ where ${\mathsf{w}(\widehat{Q})}$ is the fractional hypertree width of $\widehat{Q}$.
@@
The proof is very similar to the proof of Theorem 3.4 from Section E.1 for $\text{IVM}^{\pm}[Q]$. For every insert or delete $\delta_{\tau}\mathcal{D}=\{sR_{j}(\bm{t})\}$ where $s\in\{+,-\}$, we now have to insert an extra tuple $+\overline{R}_{j}([\tau,\tau],\bm{t})$ into $\overline{\mathcal{D}}$. This only increases the size of $\overline{\mathcal{D}}$ by a constant factor. Hence, it doesn’t affect the overall runtime.
@@
The containment $\subseteq$ in (58) holds thanks to the extra intervals $+\overline{R}_{j}([\tau,\tau],\bm{t})$ that we are inserting to $\overline{\mathcal{D}}$ in the delta version of Algorithm 2. The other containment $\supseteq$ in (58) follows from Proposition E.7. In particular, (57) only selects the leaf corresponding to $[\tau,\tau]$ in the segment tree, but Proposition E.7 tells us that this leaf could not belong to the canonical partition of any interval $[\alpha,\beta]$ unless $\tau\in\{\alpha,\beta\}$. Such intervals could only have been added to $\overline{\mathcal{D}}$ during the update $\delta_{\tau}\mathcal{D}$ because the timestamp $\tau$ is unique to this update.
@@
E.3. Upper bound in terms of $|\mathcal{D}|$ instead of $N$
@@
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$ amortized update time, where $\widehat{Q}$ is the multivariate extension of $Q$, and $|\mathcal{D}^{(\tau)}|$ is the current database size at update time $\tau$.
@@
(63) $\sum_{i\in[k-1]}|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})}=$ $\sum_{i\in[k-1]}\sum_{\tau\in[\tau_{i},\tau_{i+1})}\frac{1}{\tau_{i+1}-\tau_{i}}\cdot|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})}$
@@
E.4. Lower Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
@@
For the rest of this section, we fix a query $Q$, its time extension $\overline{Q}$ (Definition 6.2), and one component $\widehat{Q}_{\bm{\sigma}}$ in its multivariate extension $\widehat{Q}$ (Definition 2.3). Moreover, we fix an arbitrary database instance $\widehat{\mathcal{D}}_{\bm{\sigma}}$ for $\widehat{Q}_{\bm{\sigma}}$. Algorithm 3 describes our reduction from $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Theorem D.9 already gives a reduction from $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ to the (static) evaluation of $\overline{Q}$ over some database instance $\overline{\mathcal{D}}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$ that is given by Definition D.7 and satisfies the following properties:
@@
(Recall the definition of $H_{k}$ from Theorem D.9.) The above properties follow from Proposition D.8 and Theorem D.9, which also imply that $\overline{\mathcal{D}}$ can be constructed in linear time in $|\widehat{\mathcal{D}}_{\bm{\sigma}}|$, and $H_{k}(\overline{Q}(\overline{\mathcal{D}}))$ can be computed in linear time in $|\overline{Q}(\overline{\mathcal{D}})|=|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|$. It remains to show how to reduce the (static) evaluation of $\overline{Q}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q]$. This is shown by the following Lemma.
@@
  • $\overline{\mathcal{D}}\leftarrow\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$triangleright Definition D.7
@@
  • return $\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})=H_{k}(\overline{Q}(\overline{\mathcal{D}}))$triangleright Theorem D.9
Lemma E.9.
If $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $f(N)$ amortized update time for any update stream of length $N$ and some function $f$, then $\text{Eval}[\overline{Q}]$ can be solved in $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|\cdot f(2|\overline{\mathcal{D}}|)+|\overline{Q}(\overline{\mathcal{D}})|)$ time for any database $\overline{\mathcal{D}}$.
Assuming Lemma E.9 holds, Theorem 3.6 follows by taking $f(N)=N^{\omega(\widehat{Q}_{\bm{\sigma}})-1-\gamma}$ for some constant $\gamma>0$. (Recall Definition 2.2.) By Lemma E.9, this would lead into an algorithm for $\text{Eval}[\overline{Q}]$ with runtime $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\omega(\widehat{Q}_{\bm{\sigma}})-\gamma}+|\overline{Q}(\overline{\mathcal{D}})|)$, and by extension, an algorithm for $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ with runtime $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{\bm{\sigma}}|^{\omega(\widehat{Q}_{\bm{\sigma}})-\gamma}+|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|)$. This would contradict the definition of $\omega(\widehat{Q}_{\bm{\sigma}})$, thus proving that our choice of $f(N)$ is not possible. Next, we prove Lemma E.9.
Proof of Lemma E.9.
@@
Appendix F Missing Details in Section 3
@@
Proposition F.5.
For any non-hierarchical query $Q$ without self-joins, update stream of length $N$, and $\gamma>0$, there is no algorithm that solves $\text{IVM}^{\pm}[Q]$ or $\text{IVM}^{\pm}_{\delta}[Q]$ with $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ amortized update time, unless the OMv-Conjecture fails.
F.1. Proof of Proposition F.1
Proposition F.1.
@@
It follows from Proposition F.2 that the multivariate extension $\widehat{Q}$ of any non-hierarchical query $Q$ is not acyclic. This is because acyclic queries have a fractional hypertree width of $1$. It remains to prove:
Lemma F.6.
@@
We are ready to prove Lemma F.6. Consider a hierarchical query $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$ and its multivariate extension $\widehat{Q}=\bigvee_{\bm{\sigma}\in\Sigma_{k}}\widehat{Q}_{\bm{\sigma}}$ (Eq (3)). We show that each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ is acyclic. WLOG we only consider the component for $\bm{\sigma}=(1,\ldots,k)$:
@@
This query has $k-1$ atoms and contains the variables $Z_{2},\ldots,Z_{k}$ but not $Z_{1}$. Observe that $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ is the multivariate extension of a query $Q^{\prime}$ that results from $Q$ by skipping the atom $R_{1}(\bm{X}_{1})$. Hierarchical queries remain hierarchical under atom elimination. Hence, $Q^{\prime}$ is hierarchical. By induction hypothesis, $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ is acyclic. Since we obtained $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ from $\widehat{Q}_{\bm{\sigma}}$ by eliminating root variables, leaf variables, or subsumed atoms, it follows from Claim F.1 that $\widehat{Q}_{\bm{\sigma}}$ must be acyclic. This completes the induction step.
F.2. Proof of Proposition F.2
Proposition F.2.
@@
Therefore, $\rho^{*}(\widehat{Q}_{\bm{\sigma}}^{\prime})=3/2$. We use this fact to to show that $\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$ must be at least $3/2$. To that end, we use variable elimination orders (see Appendix B). In particular, we use the definition of fractional hypertree width in terms of variable elimination orders, as given by Eq. (25). Let $\bm{\mu}$ be an optimal variable elimination order for $\widehat{Q}_{\bm{\sigma}}$, i.e., $\mathsf{w}(\bm{\mu})=\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$. We make the following claim:
Claim F.2.
@@
The first equality above follows from our assumption that $\bm{\mu}$ is an optimal variable elimination order for $\widehat{Q}_{\bm{\sigma}}$. The second inequality above follows from the monotonicity of the fractional edge cover number $\rho^{*}$ (see Appendix B). ∎
F.3. Proof of Proposition F.3
Proposition F.3.
@@
Consider an arbitrary Loomis-Whitney query $Q$ of some degree $k\geq 3$ and its multivariate extension $\widehat{Q}$ with fresh variables $Z_{1},\ldots,Z_{k}$. Proposition F.2 implies that the fractional hypertree width of $\widehat{Q}$ is at least $\frac{3}{2}$. Let $\widehat{Q}_{\bm{\sigma}}$ be an arbitrary component of $\widehat{Q}$. Next, we show that the fractional hypertree width of $\widehat{Q}_{\bm{\sigma}}$ is at most $\frac{3}{2}$, which concludes the proof.
Let $\widehat{Q}_{\bm{\sigma}}^{\prime}$ be the query that we obtain from $\widehat{Q}_{\bm{\sigma}}$ by eliminating the variable $Z_{k}$. Since this variable occurs in only a single atom in $\widehat{Q}_{\bm{\sigma}}$, the fractional hypertree width of $\widehat{Q}_{\bm{\sigma}}$ and $\widehat{Q}_{\bm{\sigma}}^{\prime}$ must be the same. This follows from the equivalence between variable elimination orders and tree decompositions, as explained in Section B. It remains to show that $\widehat{Q}_{\bm{\sigma}}^{\prime}$ has a fractional hypertree width of $\frac{3}{2}$. By definition of a multivariate extension, $\widehat{Q}_{\bm{\sigma}}^{\prime}$ must contain three atoms of the form:
@@
F.4. Proof of Proposition F.4
Proposition F.4.
@@
F.5. Proof of Proposition F.5
Proposition F.5.
For any non-hierarchical query $Q$ without self-joins, update stream of length $N$, and $\gamma>0$, there is no algorithm that solves $\text{IVM}^{\pm}[Q]$ or $\text{IVM}^{\pm}_{\delta}[Q]$ with $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ amortized update time, unless the OMv-Conjecture fails.
The proof is based on the Online Vector-Matrix-Vector Multiplication (OuMv) Conjecture, which is implied by the Online Matrix-Vector Multiplication (OMv) Conjecture (Henzinger et al., 2015). See Appendix B.
Consider the simplest non-hierarchical query $Q_{nh}=R(A)\wedge S(A,B)\wedge T(B)$. The proof of Proposition F.5 consists of three parts. First, we show hardness for the problem $\text{IVM}^{\pm}[Q_{nh}]$ (Section F.5.1). Afterwards, we show how the hardness proof for $\text{IVM}^{\pm}[Q_{nh}]$ extends to a hardness proof for $\text{IVM}^{\pm}_{\delta}[Q_{nh}]$ (Section F.5.2). Finally, we explain how these two results generalize to arbitrary non-hierarchical queries without self-joins (Section F.5.3). The technique used in the latter part is standard in the literature (see, e.g. (Berkholz et al., 2017)), so we only give the high-level idea.
F.5.1. Hardness of $\text{IVM}^{\pm}[Q_{\text{nh}}]$